If f:R^n-->R is a smooth function, define hypersurface f(x)=c for some constant c, then prove velocity vector is orthogonal to the normal vector i.e <\nabla f(p),v_{\gamma}(p)>=0
1
Expert's answer
2020-10-26T19:58:12-0400

We have been provided that "f:\R^n\rightarrow \R" is "\mathscr{C}^{\infty}" . let us define the level set by


"M_f:=\{\vec x\in \R^n:f(\vec x)=c\}"

Let "\gamma:(a,b)\rightarrow \R^n" is smooth path such that "t\mapsto\gamma(t)" , for "t\in (a,b),a,b\in \R" which passes through "M_f" . let "\gamma(\theta)=p\in\R^n" for "\theta\in (a,b)" and denote "v_{\gamma}(p)=\gamma'(\theta)" is velocity vector passes through "p" .

Thus, by chain rule we get,


"\frac{d}{dt}[f(\gamma(t))]=f'(\gamma(t))v_{\gamma}(p)"

Since, "\nabla f:U\rightarrow \R^n" is smooth map from open set "U" to "\R^n" such that

"\nabla f(\vec x)=(D_1f(\vec x),...,D_nf(\vec x))"

Thus, "\nabla f(\gamma(t))^T=f'(\gamma(t))" , Hence, we get

"\frac{d}{dt}[f(\gamma(t))]=f'(\gamma(t))v_{\gamma}(p)=\frac{d}{dt}(c)=0\\
\implies <\nabla f(p),v_{\gamma}(p)>=0"

Therefore "\nabla f(p)\&v_{\gamma}(p)" orthogonal to each other.

We are done.